2/x-2+5/x+2=-6x/x^2 -4
Answer by stanbon(75887)

2/x-2+5/x+2=-6x/x^2 -4
Multiply thru by x^2-4 to get:
2(x+2) + 5(x-2) = -6x
7x-6 = 6x
x = 6
========
Cheers,
Stan H.